A 72 inches aluminum wire is to be heated from 45 f to 150 f. By how much will it expand what will be its new length?


1
Expert's answer
2019-09-05T09:13:46-0400

The linear thermal expansion is given by equation


"\Delta l=l_0\alpha\Delta T"

"=72\:\rm in\times 13\times 10^{-6}\:\rm{1/F}\times (150-45)\:\rm F=0.098\:\rm in"

The new length


"l=l_0+\Delta l=72\:\rm in+0.098\:\rm in=72.098\:\rm in"